OpenStudy (elvisiscool69):

solve.x^2-121-0 a.0 b.-11 c.11 d.11,-11

OpenStudy (solomonzelman):

\(\color{#000000 }{ \displaystyle x^2-121=0 }\) \(\color{#000000 }{ \displaystyle x^2-121\color{red}{+121}=0\color{red}{+121} }\)

OpenStudy (rose4825):

Do you understand that you have to still solve it @elvisiscool69? @SolomonZelman helped you to get started.

OpenStudy (anonymous):

11^2 = 121

OpenStudy (elvisiscool69):

i dont get it

OpenStudy (elvisiscool69):

im serious about not knowing anything about math

OpenStudy (rose4825):

Do you understand what the first post is saying?

OpenStudy (rose4825):

He moved the 0 over to the other side to start solving for x

OpenStudy (elvisiscool69):

kinda

OpenStudy (elvisiscool69):

yea i get that part

OpenStudy (rose4825):

Now you have x^2 = 121

OpenStudy (rose4825):

to get rid of the squared part you have to square-root both sides

OpenStudy (anonymous):

x^2 = 121 = 11^2 x = 11

OpenStudy (rose4825):

The second post showed you that when you square both sides you get that the square-root of 121 is 11

OpenStudy (elvisiscool69):

so 121^2=11?

OpenStudy (rose4825):

So x would equal 11

OpenStudy (elvisiscool69):

so its c

OpenStudy (rose4825):

no... \[\sqrt{x ^{2}} = \sqrt{121}\]

OpenStudy (rose4825):

x = 11

OpenStudy (rose4825):

and yes the correct answer is C

OpenStudy (elvisiscool69):

life saver!!!!

OpenStudy (elvisiscool69):

11

OpenStudy (anonymous):

Congratulations.

OpenStudy (rose4825):

@SolomonZelman has shown you the most simplest way of doing the problem. The only thing you need to understand is that to remove the 2 on top of the x, you have to square-root the x. And what you do to one side, you have to do to the other. So when you square-rooted x, you also square-rooted 121, and got your final answer; 11